OpenStudy (anonymous):

Solve for 0˚≤x≤360˚: Sin2x=sinx

Parth (parthkohli):

\[\sin(2x) - \sin(x) = 0\]\[\Rightarrow 2\sin(x)\cos(x) - \sin(x) = 0\]\[\Rightarrow \sin(x) \left(2\cos (x) - 1\right)=0\]\[\Rightarrow \cdots\]